# Coexisting phases in the chiral transition within the Linear sigma model with quarks

### v1

@@ III RESULTS AND DISCUSSION
For a fixed temperature these states form a continuum with extreme values $x_{L}$ and $x_{H}$. For a given pressure the collection of all these points, corresponding to the full range of temperatures allowed, constitutes the spinodal region. To illustrate this subject the spinodal corresponding to $P=43.5$ MeV fm-3 is shown in Fig. 2. It extends up to a maximum temperature $T=9.2$ MeV, where only states with $x_{a}\simeq x_{b}\simeq 0$ are involved. As a general trend it is found that the extension of this region reduces when i) the pressure increases, or ii) $g$ decreases, until it completely disappears.
@@
In Fig.3 the isothermal speed of sound is shown for a range of quark densities and temperatures corresponding to the first-order transition and two flavor compositions x=1/3(2/3)x=1/3\,(2/3) in the upper (lower) panel. For $x=1/3$ and $T=20$ MeV the temperature is above the CEP and $v_{S}$ has a non-monotonic trend with its minimum located approximately at the transition point (filled circle). For the lower temperatures the system experiences the coexistence of phases, which reflects in a discontinuous behavior with a sudden drop of $v_{S}$. Along this passage the speed of sound has a mild variation with the particle number, and it is approximately independent of the temperature. Furthermore, the shrinking of the spinodal region as $T$ grows is evident. A similar description applies to $x=2/3$, but in such case only the isotherms corresponding to $T=0,\text{and}\;5$ MeV undergo a first order transition.
@@
For $x=0$ (Fig.4a) the peaks nearest to $n=0$ are bounded, but they increase in magnitude when $T$ grows. On the right border, instead, there are truly divergences of the same type as those reported in Sasaki et al. (2007b, 2008) for calculations using the NJL model.
@@
For the higher $x=2/3$ a qualitative change happens. For $T=10$ MeV the system undergoes a continuous crossover and $\chi_{B}$ shows a regular behavior. The cases $T=0,\text{and}\;5$ MeV, show noticeable changes since even within the coexistence region $\chi_{B}$ has unbounded discontinuities. However it must be bear in mind that this is an extrapolation, since the effect of the weak forces will not be negligible for $x=2/3$.
The susceptibility associated with isospin is completely regular for $x=0$ (Fig.5a) with an evident enhancement for non-zero temperatures. For $x=1/3$ (Fig.5b) the unstable region is marked by peaks and drops in $\chi_{3}$, which are replaced by a monotonous trend within the coexistence of phases. Finally, for $x=2/3$ (Fig.5c) the significant deviations from the null behavior only occurs at the high-density limit of the binodal and are distinguished by a divergence for $T=5$ MeV. However the same remarks as for $\chi_{B}$, $x=2/3$ apply here.
@@ IV SUMMARY AND CONCLUSIONS
The susceptibility associated with the baryon number $\chi_{B}$, remains bounded for low asymmetry parameter $x>0$, in contrast to calculations using the unstable sector of the equation of state. As $x$ grows the divergences reappear at the high density border of the binodal, but the inclusion of the electromagnetic field could have important effects in such situations. A similar comment applies to the susceptibility $\chi_{3}$, corresponding to the isospin asymmetry.

### v2

@@ III RESULTS AND DISCUSSION
For a fixed temperature these states form a continuum with extreme values $x_{L}$ and $x_{H}$. For a given pressure the collection of all these points, corresponding to the full range of temperatures allowed, constitutes the Equilibrium Coexistence Region (CER) region. To illustrate this subject the ECR corresponding to $P=43.5$ MeV fm-3 is shown in Fig. 2. It extends up to a maximum temperature $T=9.2$ MeV, where only states with $x_{a}\simeq x_{b}\simeq 0$ are involved. As a general trend it is found that the extension of this region reduces when i) the pressure increases, or ii) $g$ decreases, until it completely disappears.
@@
In Fig.3 the isothermal speed of sound is shown for a range of quark densities and temperatures corresponding to the first-order transition and two flavor compositions x=1/3(2/3)x=1/3\,(2/3) in the upper (lower) panel. For $x=1/3$ and $T=20$ MeV the temperature is above the CEP and $v_{S}$ has a non-monotonic trend with its minimum located approximately at the transition point (filled circle). For the lower temperatures the system experiences the coexistence of phases, which reflects in a discontinuous behavior with a sudden drop of $v_{S}$. Along this passage the speed of sound has a mild variation with the particle number, and it is approximately independent of the temperature. Furthermore, the shrinking of the ECR region as $T$ grows is evident. A similar description applies to $x=2/3$, but in such case only the isotherms corresponding to $T=0,\text{and}\;5$ MeV undergo a first order transition.
@@
For $x=0$ (Fig.4a) results including the unstable region of the equation of state are shown, the peaks and drops are truly divergences of the same type as those reported and discussed in Sasaki et al. (2007b, 2008) for calculations using the NJL model.
@@
For the higher $x=2/3$ a qualitative change happens. For $T=10$ MeV the system undergoes a continuous crossover and $\chi_{B}$ shows a regular behavior. The cases $T=0,\text{and}\;5$ MeV, show noticeable changes since even within the ECR $\chi_{B}$ has large but finite discontinuities. However it must be bear in mind that this is an extrapolation, since the effect of the weak forces will not be negligible for $x=2/3$.
The susceptibility associated with isospin is completely regular for $x=0$ (Fig.5a) with an evident enhancement for non-zero temperatures. For $x=1/3$ (Fig.5b) the unstable region is marked by peaks and drops in $\chi_{3}$, which are replaced by a monotonous trend within the ECR. Finally, for $x=2/3$ (Fig.5c) the significant deviations from the null behavior only occurs at the high-density limit of the binodal and are distinguished by a divergence for $T=5$ MeV. However the same remarks as for $\chi_{B}$, $x=2/3$ apply here.
@@ IV SUMMARY AND CONCLUSIONS
The susceptibility associated with the baryon number $\chi_{B}$, remains bounded for low asymmetry parameter $x>0$, in contrast to calculations using the unstable sector of the equation of state. As $x$ grows large but finite discontinuities appear at the high density border of ECR, but the inclusion of the electromagnetic field could have important effects in such situations. A similar comment applies to the susceptibility $\chi_{3}$, corresponding to the isospin asymmetry.